OpenStudy (anonymous):

For what value of x is the rational expression below equal to zero? x-4/x-6

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ \normalsize \text{the expression is undefined when it is infinity} \hspace{.33em}\\~\\ \dfrac{2x+4}{x-7}=\infty \hspace{.33em}\\~\\ \dfrac{2x+4}{x-7}=\dfrac{1}{0} \hspace{.33em}\\~\\ \dfrac{x-7}{2x+4}=\dfrac{0}{1} \hspace{.33em}\\~\\ x-7=0 \hspace{.33em}\\~\\ \Large \fbox{x=7} \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

@mathmath313, I don't understand your answer. For what value is the rational expression equal to zero? Why make it equal to infinity? Why then make it to division by zero, which of course is undefined. and finally invert this. Surely the question being asked is: 2x +4/x-7 = 0 ?

OpenStudy (mathmath333):

wait he changed the question lol it was different earlier

OpenStudy (mathmath333):

brittany if u want to ask another question , u can close this question and ask in a new post ,thanks

OpenStudy (anonymous):

you're fine @mathmath333

OpenStudy (anonymous):

yea I know don't have time to

OpenStudy (mathmath333):

ok lol

OpenStudy (anonymous):

firstly the question is x-4/x-6 not 2x+4/ x-7

OpenStudy (anonymous):

we all make mistakes! and this should not reflect on @mathmath333's ability.

OpenStudy (mathmath333):

lol i didnt made any mistake

OpenStudy (anonymous):

sorry, I thought the question asked was x-4/x-6?

OpenStudy (anonymous):

He didn't make a mistake I just refused to close the question so I changed it because I'm on a time limit

OpenStudy (anonymous):

didnt know that!

OpenStudy (anonymous):

can someone help me please with the new question

OpenStudy (anonymous):

yes of course, both me and @mathmath333 can help you

OpenStudy (anonymous):

Thank You

OpenStudy (mathmath333):

did u solved this (x-4/x-6)

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ \dfrac{x-4}{x-6}=0 \hspace{.33em}\\~\\ x-4=0 \hspace{.33em}\\~\\ x=4 \hspace{.33em}\\~\\ }\end{align}\)